If 4 fifth X _____ = 2 fifth What is _____?
Answer by joey_quizan(15)

solution:
4/5 x _____ = 2/5
let y be the missing number
4/5 x y = 2/5 ----devide both sides of the equation by 4/5
y = 2/5 / 4/5
y = 2/5 * 5/4
y = 1/2 --- answere